# Planning from Imagination: Episodic Simulation and Episodic Memory for Vision-and-Language Navigation

### v1

@@ Abstract
Humans navigate unfamiliar environments using the capabilities of episodic simulation and episodic memory. Developing imagination-based memory, analogous to episodic simulation and episodic memory, can enhance embodied agents’ comprehension of the complex relationship between environments and objects. However, existing Vision-and-Language Navigation (VLN) agents fail to perform the aforementioned mechanism. We propose a novel architecture to help agents build a recurrent imaginative memory system. Specifically, the agent can maintain a reality-imagination hybrid global memory during navigation and expand the memory map through imaginative mechanisms and navigation actions. Correspondingly, we design a series of pre-training tasks to help the agent acquire fine-grained imaginative abilities. Our agents improve the state-of-the-art (SoTA) success rate (SR) by 7% while simultaneously imagining high-fidelity RGB representations for future scenes.
@@ 1 Introduction
Autonomous embodied agent navigation is advancing through Vision-and-Language Navigation (VLN) research (Anderson et al. 2018; Qi et al. 2020). In a VLN task, the agent needs to reach the target location given navigation instructions (e.g. “Exit the room then go straight and turn left. Go straight until you pass a TV.”). This task, however, is more challenging when agents navigate in unseen environments. The navigation performance is significantly degraded compared to that in the seen environment.
@@
Thus, it is natural to focus on human navigation behaviors in unseen environments, and two important neuroscience concepts were identified: the episodic simulation and the episodic memory (Kühn and Gallinat 2014; Gomez, Rousset, and Baciu 2009), which means using memories to imagine uncertain information or scenes, aiding future decision-making mentally). According to their navigation memory, humans can imagine both fine-grained knowledge (e.g., RGB images) and higher-level spatial information (e.g., location distribution) for the future in unseen environments (Tang, Yan, and Tan 2017). For example, if the navigation instruction is “go to the laundry room” (Fig 1.), humans would imagine a room with washing machines, likely to be adjacent to a toilet they passed earlier. Such a simulation would provide clues for navigating directions.
@@
Existing approaches equip agents with mechanisms to infer or imagine unseen environments. However, these agents lack the ability to integrate their imaginative outputs into long-term memory, preventing the formation of episodic simulations and persistent memories (Wang et al. 2023a; An et al. 2023). While current methods enable agents to predict potential object locations (Shah et al. 2023; An et al. 2024) and RGB features (Moudgil et al. 2021; Qi et al. 2021), such predictions are often ephemeral and transient, overwritten at each navigation step. This transient nature prevents agents from constructing durable memory representations that fuse visual and contextual information to support future reasoning or decision-making. For instance, in Fig 1, the inability to infer the presence of spice bottles from contextual cues such as a knife and fork highlights the limitations of existing imagination mechanisms. Consequently, these methods produce isolated and short-lived imaginative outputs, falling short of the continuity and compositional depth characteristic of human episodic memory and simulation.
@@
Figure 1: Humans utilize episodic memory and episodic simulation to recall past experiences and predict future outcomes in unfamiliar environments. However, navigation agents often struggle in unseen environments due to their inability to construct and leverage such cognitive frameworks.
@@
In this paper, we introduce the Space-Aware Long-term Imaginer (SALI) agent, the first agent designed to emulate human-like episodic simulation and memory for navigation. The SALI agent leverages an imagination-based memory system to capture both high-level spatial information about the navigation environment and fine-grained local features. By incorporating commonsense knowledge (e.g., associating specific objects with corresponding rooms, such as a refrigerator with a kitchen), the SALI agent builds and maintains a long-term imagination-based memory, as illustrated in Fig 2. This memory structure enables the agent to integrate imaginative outputs with stored information to support informed action decisions. To implement this capability, we model the agent’s memory using a topological map structure (Chen et al. 2022), where nodes represent spatial locations and are enriched with local RGBD data and semantic information generated by the imagination component. Multimodal transformers are employed to encode these nodes alongside natural language instructions, enabling the agent to make effective navigation decisions at each step. The combination of imagination and memory ensures that the SALI agent can reason holistically about its environment while adapting dynamically to new scenarios. As a result, the SALI agent achieves state-of-the-art performance on two widely recognized VLN benchmarks, demonstrating the effectiveness of our approach.
@@
  • Our proposed SALI agent has human-like episodic memory and episodic simulation abilities, enhancing navigation performance in unseen environments. Viewing from the perspective of human brains, we endow the agent with the ability to learn commonsense knowledge through imagination and memory.
  • We established a recurrent and end-to-end imagination mechanism that generates high-fidelity RGB scene representations. The SALI agent is capable of imagining realistic scene images and dynamically fusing the imagination into a real-imaginary combined memory. The agent can make navigation decisions based on this hybrid map, developing more effective navigation skills.
  • The SALI agent achieves state-of-the-art (SoTA) performance on VLN benchmarks. In the unseen environments, it improves 5% SR on the R2R dataset and 7% on the REVERIE dataset.
@@ 2 Related Works
Vision-and-language Navigation. Using visual inputs and natural language to guide robot navigation is a fundamental task in embodied AI (Anderson et al. 2018; Qi et al. 2020), where effectively aligning multimodal visual-linguistic inputs remains a core challenge (Lin et al. 2022). Existing VLN approaches (Lu et al. 2019; Qi et al. 2021; Hong et al. 2021) often struggle in unseen environments, where agents fail to correlate visual observations, particularly unfamiliar materials and textures, with navigation instructions, resulting in degraded performance. To address this limitation, we propose a novel VLN architecture inspired by human episodic simulation and memory, enabling agents to enhance navigation robustness and performance in unseen environments.
@@
Memory Mechanism. Constructing memory to provide long-term decision support is important for embodied agents. SLAM has been widely used to construct memory maps, especially in navigation (Chaplot et al. 2020; Temeltas and Kayak 2008). To reduce memory overhead and enable more efficiency, the construction of memory mechanisms based on 2D bird-eye-view maps (An et al. 2023; Liu et al. 2023) or ego-centric grid maps (Georgakis et al. 2022; Wang et al. 2023c) can also help agents understand spatial information. Additionally, some navigation work uses topological maps to further simplify and extract high-level features to help the robot better learn spatial structure information (An et al. 2023). However, the correlation between historical memory and future imagination remains largely unexplored, leaving the potential of integrating the two underutilized. This oversight reduces the robustness of navigation agents and significantly impairs their performance in unseen environments. Meanwhile, the construction of abstract anticipatory and foresight thinking is necessary for the construction of human-like robots. We constructed a memory mechanism analogous to human episodic simulation memory to acquire knowledge about future scenarios.
@@
Prediction during Navigation. Inspired by video prediction (Mallya et al. 2020), prior works have adopted imagery mechanisms to support navigation decisions, including predicting future spatial structures from depth camera data (Shah et al. 2023; Cui et al. 2019) and generating fine-grained visual information. They use RGB images to generate high-fidelity images (Moudgil et al. 2021; Wang et al. 2023a). Furthermore, there is also pioneering work that explores the generation of both spatial information and image information (Wang et al. 2018b; Li et al. 2023). Instead of simply applying imagination mechanisms in isolation, we leverage a memory-based framework to fully realize the potential of imaginative predictions. By integrating imagination into a memory mechanism, we enable the generation and maintenance of long-term, reusable navigation information. To achieve this, we propose an end-to-end adaptive image maintenance and generation mechanism that bridges the gap between transient predictions and persistent memory.
@@
Figure 2: We propose building a hybrid imagination-reality memory for long-term navigation decisions. Based on the navigation observation and encoded features, the agent will generate a local imagination tree. This tree will then be fused into a hybrid memory, which the agent can utilize to make navigation decisions. The figure also illustrates a series of pre-training tasks that we propose for the imagery module.
@@ 3 Method
Problem Formulation. Our problem is set in a discrete VLN environment $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ (Anderson et al. 2018), where $\mathcal{V}$ represents navigation nodes and $\mathcal{E}$ represents the connectivity relations. The VLN dataset provides the instruction set $I$ and the corresponding ground-truth navigation paths $\hat{P}$. The VLN task requires guiding a robot to accomplish navigation tasks under $I$. At each navigation step $t$, the agent can acquire the RGB image $r_{t}$ and depth image $d_{t}$ via the camera and the current position information $p_{t}$ via the GPS sensor. Following (Irshad et al. 2022), agents are equipped with a visual classifier to acquire semantic images $s_{t}$. with observation $O_{t}=\{r_{t},d_{t},s_{t},p_{t}\}$, the agent needs to learn a policy $\pi(I,O_{t})$ to decide the current action $a_{t}$.
Method Overview. As shown in Fig 2, our SALI agent has a human-like episodic simulation and episodic memory mechanism. At each navigation step $t$, the agent will maintain a topological map as its memory to store both realistic and imagined information and make navigation decisions based solely on the map (Section 3.1). Then, the agent will use its memory to imagine future information of both high-level spatial knowledge and low-level image features and merge the imagery into the memory (Section 3.2). We conclude this part by presenting our approach to training and inference in Section 3.3.
Figure 3: The imagination model includes four pre-trained models (inpaint model, spade model, room-type model, and waypoint model). We propose an end-to-end architecture that allows local imagery trees to continuously update and maintain themselves, producing high-quality depth images, semantic images, and RGB images
@@ 3.1 Real-Imaginary Hybrid Memory
To provide the agent with the possibility of global action planning, we construct a mixed-granularity map memory structure. As shown in Fig 2, at navigation step $t$, the agent will update a topological map $G_{t}$ based on the current observation $O_{t}$ and historical information $H_{t}$.
@@ 3.1.1 Memory Map Representation
The global map can be represented as $G_{t}=\{N_{t},E_{t}\}$. Where the $E_{t}$ matrix records the Euclidean distance between neighboring nodes. $N_{t}$ represents the nodes in the map and records $V_{t}$, $p_{t}$ and $f_{t}$ information. $V_{t}=\left\{r_{t},d_{t},s_{t}\in\mathbb{R}^{H\times W\times C}\right\}$ in equirectangular panorama form will be used to generate the fine-grained images. Furthermore, we use the pre-trained Vision Transformer (ViT) (Lu et al. 2019) and ResNet encoder to get the image feature $f_{t}$, which is also stored in node $N_{t}$.
Stored nodes can be classified into four categories: visited nodes [Uncaptioned image], current nodes [Uncaptioned image], navigable nodes [Uncaptioned image], and imagination nodes [Uncaptioned image]. The information stored in these four types of nodes varies. For $V_{t}$, the complete and fixed image information has been stored in the visited node and current node. The navigable node’s image feature is represented by partial image features observed from neighboring real nodes and may contain information from imagination nodes. The $V_{t}$ of the imagination node is completely represented by the imagined image. For $p_{t}$, the agent cannot reach location $p_{t}$ of [Uncaptioned image], [Uncaptioned image], or [Uncaptioned image] via action $a_{t}$ (Anderson et al. 2018). Only the navigable node’s $p_{t}$ is the potential target location. In the map $G_{t}$, global action space is $A^{G_{t}}=\{a_{t}^{1},...a_{t}^{k}\}$, where $k$ is the number of navigable nodes in the map.
To reduce the memory load, we perform a pruning operation on the newly imagined nodes. The criteria of pruning are based on the sum of the feature cosine similarity and the negative position mean square error (MSE) between the imagination node $N_{i}$ and other navigable/imagination node $N_{j}$. If it is the same node, the node attributes are updated with an average pooling of the images and features from both nodes. The new node remains an imagination node when both are imagination nodes, otherwise is labeled as a navigable node. In addition, we set an upper bound of $\bar{N}$ on the number of imagination nodes to reduce memory load.
$\text{Criterion}(N_{i},N_{j})=\frac{f_{i}f_{j}}{||f_{i}||||f_{j}||}$ $-\text{MSE}(p_{i},p_{j})$ (1)
@@ 3.1.2 Dynamic Action Planning
As shown in Fig 2, we then input the topological memory $G_{t}$ at time $t$ into a multimodal transformer to obtain representations. this multimodal transformer contains pre-trained real node ( [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] ) encoders, fine-tuning imagination node ( [Uncaptioned image] ) encoders, and a pre-trained instruction encoder. The real node encoder and the imagination node encoder have the same structure and are therefore unified as node encoder for elaboration.
Node Embedding. The node visual feature $\hat{V}_{t}$ comprises three elements: node feature $f_{t}$, location encoding, and navigation step encoding. The latter marks visited nodes with their last visited time (00 for [Uncaptioned image], [Uncaptioned image]). This encoding reflects the temporal structure of memory. $\hat{V}_{t}$ is categorized into $\hat{V}_{t}^{r}$ and $\hat{V}_{t}^{i}$ based on whether the nodes are real or imaginary. Furthermore, a ‘stop’ node is added to the memory to represent a stop action and is connected to all other nodes.
Instruction and Node Encoders. Each word embedding in $I$ is augmented with a positional embedding and a type embedding. A multi-layer transformer processes these tokens to generate contextual representations, denoted as $\hat{I}$
Node and word embeddings are fed into a multi-layer cross-modal transformer. Inspired by (Chen et al. 2022), we endow it with a graph-aware self-attention (GASA) layer to capture the environment layout. The network output is navigation scores $s_{i}^{r}$ for navigable nodes and $s_{i}^{i}$ for imagination nodes. FFN below is a two-layer feed-forward network.
$s_{t}^{k}=\text{FFN}(\text{GASA}(\hat{V}_{t}^{k})),k=i,r$ (2)
Action Fusion Policy. Our work introduces an adaptive framework for navigation decision-making, which involves the dynamic integration of navigation scores obtained from both navigable nodes and imagination nodes. Initially, navigation scores from imagination nodes will be multiplied by a fusion factor and added to the scores of the nearest navigable nodes in Euclidean distance, since they are not reachable. The fusion factor is generated by concatenating the imagination nodes and real nodes and then feeding it into an FFN layer. The network structure is illustrated as follows.
$\gamma_{t}=\text{Sigmoid}(\text{FFN}(\left[\hat{V}^{r}_{t},\hat{V}^{i}_{t}\right]))$ (3)
$\hat{s_{t}}=s_{t}^{r}+\sum\nolimits_{j\in\mathcal{S}(i)}\gamma_{t}s^{i}_{t,j}$ (4)

### v2

@@ Abstract
Humans navigate unfamiliar environments using episodic simulation and episodic memory, which facilitate a deeper understanding of the complex relationships between environments and objects. Developing an imaginative memory system inspired by human mechanisms can enhance the navigation performance of embodied agents in unseen environments. However, existing Vision-and-Language Navigation (VLN) agents lack a memory mechanism of this kind. To address this, we propose a novel architecture that equips agents with a reality-imagination hybrid memory system. This system enables agents to maintain and expand their memory through both imaginative mechanisms and navigation actions. Additionally, we design tailored pre-training tasks to develop the agent’s imaginative capabilities. Our agent can imagine high-fidelity RGB images for future scenes, achieving state-of-the-art result in Success rate weighted by Path Length (SPL).
@@ 1 Introduction
Autonomous embodied agent navigation is advancing through Vision-and-Language Navigation (VLN) research (Anderson et al. 2018b; Qi et al. 2020). In VLN tasks, an agent needs to reach the target location given navigation instructions. This task, however, is more challenging when agents navigate in unseen environments, with performance degradation compared to seen environments.
@@
Human navigation behaviors in unfamiliar environments provide valuable insights, particularly through the neuroscience concepts of episodic simulation and memory (Kühn and Gallinat 2014; Gomez, Rousset, and Baciu 2009), which may help address this challenge. These mechanisms enable humans to use memory to mentally simulate uncertain information or scenes, aiding decision-making. Humans can imagine fine-grained visual features (e.g., RGB images) and higher-level spatial structures (e.g., location distributions) in unseen environments (Tang, Yan, and Tan 2017). As illustrated in Figure 1, given an instruction like “go to the laundry room”, humans might imagine a room with washing machines, likely adjacent to a bathroom encountered earlier, using such mental simulations to infer navigation directions.
@@
While existing approaches have incorporated mechanisms for imagining unseen environments, they lack the ability to integrate the imagined results into time-series persistent memory. Specifically, This prevents the formation of long-term episodic simulations, which is essential for effective navigation (Wang et al. 2023a; An et al. 2023). Current methods enable agents to predict potential object locations (An et al. 2024) and RGB features (Moudgil et al. 2021; Qi et al. 2021), but these predictions are transient and overwritten at each step, precluding durable memory representations. Such limitations hinder agents from fusing visual and contextual information to support future reasoning or decision-making. For instance, in Figure 1, an agent’s inability to infer the presence of spice bottles from contextual cues such as a knife and fork illustrates the shortcomings of existing mechanisms. Consequently, these methods produce isolated and short-lived imaginative outputs, lacking the continuity and compositional depth characteristic of human episodic memory and simulation.
@@
Figure 1: Humans utilize episodic memory and episodic simulation to recall past experiences and predict future outcomes in unfamiliar environments. In contrast, navigation agents often struggle in unseen environments due to their inability to construct and leverage such cognitive frameworks.
@@
In this paper, we introduce the Space-Aware Long-term Imaginer (SALI) agent, the first navigation agent explicitly designed to emulate human-like episodic simulation and memory. SALI leverages an imaginative memory system to capture both high-level spatial structures of the environment and fine-grained RGB features. The generation of these imaginative outputs is achieved by integrating prior imagined results with historical information, forming a recurrent imagination module. This iterative process enables SALI to maintain a hybrid memory system that combines real and imagined representations, facilitating robust reasoning and navigation in complex and unseen environments. To implement this system, we model the agent’s memory using a topological map (Chen et al. 2022), where nodes represent spatial locations enriched with RGB-D and semantic features generated by the imagination module or the navigation actions. Multimodal transformers then encode these nodes alongside natural language instructions, enabling reasonable navigation decisions at each step. The integration of imaginative and realistic information allows SALI to reason holistically about its environment while dynamically adapting to new scenarios. As a result, SALI achieves state-of-the-art (SoTA) performance on R2R (Anderson et al. 2018b) and REVERIE (Qi et al. 2020), highlighting its effectiveness.
@@
  • Our proposed SALI has human-like episodic memory and episodic simulation abilities, enhancing general navigation performance. Viewing from the perspective of human brains, we endow the agent with the ability to learn anticipatory knowledge through imagination.
  • We established a recurrent, end-to-end imagination module that generates high-fidelity RGB scene representations. SALI dynamically fuses imagined scenes with real observations into a hybrid memory map, ensuring effective navigation decisions.
  • SALI achieves SoTA performance on R2R and REVERIE, improving SPL by 8% and 4% in unseen scenarios, respectively.
@@ 2 Related Works
Vision-and-language Navigation. Guiding robot navigation using visual inputs and natural language instructions is a core task in embodied AI (Anderson et al. 2018b; Qi et al. 2020). A key challenge lies in effectively aligning multimodal visual-linguistic inputs (Lin et al. 2022). Existing VLN approaches (Lu et al. 2019; Qi et al. 2021; Hong et al. 2021) often falter in unseen environments, struggling to associate unfamiliar visual observations, such as novel materials and textures, with navigation instructions, leading to degraded performance. To address these limitations, we propose a novel VLN architecture inspired by human episodic simulation and memory, enhancing robustness and performance in unseen environments.
@@
Memory Mechanism. Constructing memory for long-term decision support is critical for embodied agents. SLAM has been widely used to build memory maps in navigation (Chaplot et al. 2020; Temeltas and Kayak 2008). To reduce memory overhead, more methods such as 2D bird-eye-view maps (An et al. 2023; Liu et al. 2023) or ego-centric grid maps (Georgakis et al. 2022; Wang et al. 2023c) are investigated to help agents better understand spatial information. Additionally, some other navigation agents employ topological maps to further simplify and extract high-level features to learn underlying spatial knowledge (An et al. 2023). However, Integrating historical memory with future imagination is underexplored, limiting navigation robustness in unseen environments. Drawing inspiration from human cognition, we design an episodic simulation-based memory mechanism to enhance agents’ foresight.
@@
Prediction during Navigation. Inspired by video prediction (Mallya et al. 2020), prior works have adopted imaginary mechanisms to support navigation decisions by generating fine-grained visual information. They use RGB-D images to generate high-fidelity images (Moudgil et al. 2021; Wang et al. 2023a). Furthermore, there is also pioneering work that explores the generation of spatial structural information (Wang et al. 2018b; Li et al. 2023) with the help of depth cameras (Shah et al. 2023; Cui et al. 2019). However, xisting imagination mechanisms operate in isolation, without integration into persistent memory. We propose a memory-based framework to bridge transient predictions with long-term memory, enabling reusable navigation information. Our adaptive mechanism integrates imagination into memory for durable, context-aware navigation.
@@
Figure 2: We propose building a hybrid imagination-reality memory for long-term navigation decisions. Based on the navigation observation and trajectories, the agent will imagine future scenes for unvisited environments. The imagination will then be fused into its hybrid memory to aid further decision-making. The figure also illustrates a series of pre-training tasks that we propose for the imagination module.
@@ 3 Method
Problem Formulation. The VLN task is set in a discrete environment $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ (Anderson et al. 2018b), where $\mathcal{V}$ is navigation nodes and $\mathcal{E}$ represents the connectivity relations. The VLN dataset provides instruction set $I$ and corresponding ground-truth navigation paths. It requires guiding a robot to accomplish navigation tasks with given instructions. At each navigation step $t$, the agent can acquire the RGB image $r_{t}$ and depth image $d_{t}$ via the camera and the current position information $p_{t}$ via the GPS sensor. Following (Irshad et al. 2022), an agents is equipped with a visual classifier to acquire semantic images $s_{t}$. The agent needs to learn a policy to make actions based on observations $O_{t}=\{r_{t},d_{t},s_{t},p_{t}\}$ and instruction.
Method Overview. As shown in Figure 2, SALI has a human-like episodic simulation and episodic memory mechanism. At each navigation step $t$, the agent will maintain a topological map as its memory to store both realistic and imaginative information and make navigation decisions based solely on the memory (Section 3.1). Then, the agent will use its memory to imagine future information of both high-level spatial knowledge and low-level image features and merge the imaginary into the memory (Section 3.2). We conclude this part by presenting our approach to training and inference in Section 3.3.
@@ 3.1 Real-Imaginary Hybrid Memory
To provide the agent with the ability of global action planning, we construct a mixed-granularity memory. As shown in Figure 2, at navigation step $t$, the agent will update the topological memory $G_{t}$ based on the current observation $O_{t}$ and historical information.
@@ 3.1.1 Memory Map Representation
The topological map memory is represented as $G_{t}=\{N_{t},E_{t}\}$, where $E_{t}$ records the Euclidean distance between neighboring nodes, and $N_{t}$ represents the nodes containing visual inputs $V_{t}$, position information $p_{t}$, and node feature $f_{t}$. Visual inputs $V_{t}=\left\{r_{t},d_{t},s_{t}\right\}$, stored in equirectangular panorama form, is used to generate fine-grained images. We use a pre-trained Vision Transformer (ViT) (Lu et al. 2019) and ResNet encoder to get the node feature $f_{t}$ from visual inputs $V_{t}$.
Nodes $N_{t}$ can be classified into four categories: visited nodes [Uncaptioned image], current nodes [Uncaptioned image], navigable nodes [Uncaptioned image], and imagination nodes [Uncaptioned image]. Each node type stores different information. For $V_{t}$, visited and current nodes store complete and fixed image information, while the navigable node’s $V_{t}$ is represented by partial visual inputs observed from neighboring real nodes and contains potential information from imagination nodes. The $V_{t}$ of the imagination node is completely generated from imagined images. For $p_{t}$, the agent can only reach the potential position associated with navigable nodes, as locations corresponding to [Uncaptioned image], [Uncaptioned image], or [Uncaptioned image] are not accessible via action $a_{t}$.
To manage memory efficiency, pruning operations are applied to newly imagined nodes. If two nodes are determined to represent the same node, the pruning operation retains a single node, and its feature $f_{t}$ is updated using the average pooling of the features from both nodes. The new node remains an imagination node when both are imagination nodes, otherwise is labeled as a navigable node. In addition, we set an upper bound of $\bar{N}$ on the number of imagination nodes. The criteria of pruning are based on the feature cosine similarity and the negative position mean square error (MSE) between the imagination node $N_{i}$ and other node $N_{j}$:
$\text{Criterion}(N_{i},N_{j})=\frac{f_{i}f_{j}}{||f_{i}||||f_{j}||}$ $-\text{MSE}(p_{i},p_{j}).$ (1)
Figure 3: The imagination model includes four pre-trained models (inpaint model, spade model, room-type model, and waypoint model). We propose an end-to-end architecture that allows local imaginary trees to continuously update and maintain themselves, producing high-quality depth images, semantic images, and RGB images.
@@ 3.1.2 Dynamic Action Planning
As shown in Figure 2, the topological memory $G_{t}$ at time $t$ is processed by a multimodal transformer to obtain contextual representations. this transformer contains pre-trained encoders for real nodes ( [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] ), fine-tuned encoders for imagination nodes ( [Uncaptioned image] ), and a pre-trained instruction encoder. The real node encoder and the imagination node encoder share the same structure, unified as the node encoder for simplicity.
Node Embedding. The input to generate node embeddings $\hat{V}_{t}$ includes three elements: node feature $f_{t}$, location encoding, and navigation step encoding. The latter marks visited nodes with their last visited time (00 for [Uncaptioned image], [Uncaptioned image]). This encoding reflects the temporal structure of memory. Node embeddings $\hat{V}_{t}$ are categorized into $\hat{V}_{t}^{r}$ and $\hat{V}_{t}^{i}$ based on whether the nodes are real or imagined. Furthermore, a ‘stop’ node is added to the memory to represent a stop action and is connected to all other nodes.
Instruction and Node Encoders. Each word embedding in instruction $I$ is augmented with a positional embedding and a type embedding. A multi-layer transformer processes these tokens to generate contextual representations, denoted as instruction embeddings $\hat{I}$.
Node and word embeddings $\hat{V}_{t}$ and $\hat{I}$ are fed into a multi-layer cross-modal transformer. Following (Chen et al. 2022), we endow the transformer with a graph-aware self-attention (GASA) layer to capture the environment layout. The network outputs navigation scores $s_{i}^{r}$ and $s_{i}^{i}$ for each navigable and imagination node as follows, where FFN is a two-layer feed-forward network:
$s_{t}^{k}=\text{FFN}(\text{GASA}(\hat{V}_{t}^{k})),k=i,r.$ (2)
Action Fusion Policy. We propose an adaptive framework for navigation decision-making, which involves the dynamic integration of navigation scores obtained from both navigable nodes and imagination nodes. Initially, navigation scores from imagination nodes will be multiplied by a fusion factor $\gamma_{t}$ and added to the scores of the nearest navigable nodes in Euclidean distance, since they are not reachable. The fusion factor is generated by concatenating the imagination nodes and real nodes and then feeding it into an FFN layer. The scores for next-node selection are obtained by:
$\gamma_{t}=\text{Sigmoid}(\text{FFN}(\left[\hat{V}^{r}_{t},\hat{V}^{i}_{t}\right])),$ (3)
$\hat{s_{t}}=s_{t}^{r}+\sum\nolimits_{s^{i}_{t}\in\mathcal{S}(i)}\gamma_{t}s^{i}_{t},$ (4)